Rieger International is attempting to evaluate the feasibility of investing $93,000 in a piece of equipment that has a 5​-year life. The firm has estimated the cash inflows associated with the proposal as: The firm has a 12% cost of capital.

1   $30,000
2   $25,000
3   $30,000
4   $35,000
5   $35,000


a.  Calculate the payback period for the proposed investment in years and months

b.  Calculate the net present value (NPV) for the proposed investment.

c.  Calculate the internal rate of return (IRR)​, rounded to the nearest whole​ percent, for the proposed investment.

d.  Evaluate the acceptability of the proposed investment using NPV and IRR. What recommendation would you make relative to implementation of the​ project
Answer #1
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate=Cost of Capital =12%=0.12
N=Year   of Cash Flow
CASH FLOW ANALYSIS OF EXPANSION PROJECT
N Year 0 1 2 3 4 5
A Total initial cash flow -$93,000.00
B Annual Cash In Flow $30,000 $25,000 $30,000 $35,000 $35,000
C=A+B Net Cash Flow ($93,000) $30,000 $25,000 $30,000 $35,000 $35,000
Cumulative Cash Flow ($93,000) ($63,000) ($38,000) ($8,000) $27,000 $62,000 SUM
PV=C/(1.12^N) Present Value of Net Cash Flow ($93,000) $26,786 $19,930 $21,353 $22,243 $19,860 $17,172
b NPV=Sum of PV Net Present Value(NPV) $17,172
a Payback Period =Period at which Cumulative cash flow=NIL
Payback Period =3+(8000/35000)                     3.23 Years
Payback Period=3 years 2.74 months
c Internal Rate of Return 19% (Using IRR function of excel over the Net Cash Flow)
d NPV is positive
IRR is greater than Cost of Capital
Hence the investment is acceptable in both criterion